# Can femtoscopic correlation function shed light on the nature of the lightest, charm, axial mesons?

### v1

@@ III.2 Lednicky-Lyuboshits approximation
To shed some light on the interpretation of the CFs, it can be instructive to review the Lednicky-Lyuboshits (LL) model, which is based on replacing the full wave function for a single channel by its non-relativistic, asymptotic ($r\to\infty$) form, corresponding to the superposition of plane and converging spherical waves Lednicky and Lyuboshits (1981). In particular, we benefit from the discussion presented in the Appendix of Ref. Albaladejo et al. (2023) and Sections V.B and V.C of Ref. Kamiya et al. (2022), which have some of their fundamental aspects reproduced here.
Proceeding ahead, the consideration of the LL approximation, together with a Gaussian source, and using the relationship between the standard quantum mechanics amplitude $f(k)$ and the scattering matrix $T$, i.e. $f(k)=-T/(8\pi\sqrt{s})$, allow us to write the single-channel CF as Fabbietti et al. (2021); Lednicky and Lyuboshits (1981); Kamiya et al. (2022); Albaladejo et al. (2023)
@@
Thus, considering an attractive interaction generated by the strong force, the CF given in Eqs. (29) and (30) behaves as follows. Near the threshold, for negative values of the scattering length (which means an unbound scenario for the system) the CF acquires (i) a strong enhancement when $|a|\gg R$ (i.e. smaller source); (ii) a moderate enhancement when $|a|\sim R$; and (iii) a value $C_{LL}(0)\gtrsim 1$ when $|a|\ll R$ (larger source). On the other hand, in the situation of $a>0$, corresponding to an attractive interaction which could generate a bound or quasi-bound state, in the low-momentum limit, the CF achieves (i) a strong enhancement for smaller sources $(a\gg R)$; (ii) a value very close to one when $a/R\sim 4/\sqrt{\pi}\sim 2.3$; (iii) its minimum value ($\simeq 0.4$) when $a/R\sim 2/\sqrt{\pi}\simeq 1.1$; (iv) a moderate dip when $a<R$; and (v) a value $C_{LL}(0)\lesssim 1$ for larger sources ($a\ll R$). This behavior is summarized in Fig. 7 (left panel), where one can see that the enhancement of CF at a given value of $R$ is not conclusive concerning the formation of a bound or quasi-bound state. Notwithstanding, as a consequence of the dependence of the CF on $R$ and $a$, one can infer the existence of a bound or quasi-bound state when, near the threshold, the CF moves from an enhancement to a dip at $a\simeq R$ as $R$ increases. In this sense, experimental analyses of the CF in systems with different sizes, for instance, $pp,pA$, and $AA$ collisions, deserve special attention.
@@
Then, considering that a resonance present at $k_{R}$ generates $\delta(k_{R})=\pi/2$, which when used in Eq. (31) yields $C_{LL}(x_{R})=1+\frac{e^{-4x_{R}^{2}}}{2x_{R}^{2}};\ x_{R}=k_{R}R$. Accordingly, the CF ends up having the following properties (i) $C_{LL}^{\prime}(x_{R})<0$; (ii) $C_{LL}<1$ for $x\gtrsim x_{R}$; and (iii) $C_{LL}\simeq 1$ for large $x_{R}$. To be more didactic, from the analytical expression for the Breit-Wigner-like phase shift, $\delta(E)=\Gamma(E)/(2E_{R}-2E)$ ($\Gamma$ being the width and $E_{R}$ the energy at $k_{R}$), one gets $\cot{\delta(k)}\sim-(x^{2}-x_{R}^{2})/x_{\Gamma}^{2}$, with $x_{\Gamma}^{2}=\mu\Gamma R^{2}$ ($\mu$ being the reduced mass of the two particles in the channel). Thus, the use of this last expression of $\cot{\delta(k)}$ in Eq. (31) can engender in $C_{LL}$ (i) a maximum at $k\lesssim k_{R}$ and a pronounced minimum at $k\gtrsim k_{R}$ for small $x_{\Gamma}$, (ii) a weakened minimum at $k\gtrsim k_{R}$ for intermediate values of $x_{\Gamma}$, and (iii) an almost plateau-like appearance at $k\lesssim k_{R}$ for large $x_{\Gamma}$. This behavior is summarized in Fig. 7 (right panel), in accordance with that in Ref. Albaladejo et al. (2023).
@@ III.3.1 CFs in isospin basis
Figure 8: CFs for the $D^{*}\pi$ (top panels) and $D\rho$ (bottom panels) channels as functions of their relative momentum $k$, taking different values of the source size. The left, center, and right panels show the results obtained respectively with $I=1/2$ considering the model A, with $I=1/2$ considering model B, and with $I=3/2$. Recall that the interaction in the isospin 3/2 configuration is repulsive and weaker as compared to the case of isospin 1/2. Thus, no states are formed in the isospin 3/2 case and the amplitudes are obtained from the same model [using Eq. (2)].
@@
In Fig. 8 we plot the results of the correlation functions for the most relevant channels $D^{*}\pi$ and $D\rho$ as functions of their CM momentum $k$, for different values of the range parameter $R$ of the source, considering the following scenarios: $I=1/2$ with the model A, $I=1/2$ with the model B and $I=3/2$. For the sake of comparison and to reach a more profound comprehension concerning our findings, the results with the single-channel LL approximation are also included.
@@
First, one can notice the distinct behavior of the $C_{D^{*}\pi}^{(1/2)}(k)$ when models A or B are employed ( see left and center panels in the upper row of Fig. 8). To start with the discussions, let us focus on the curves related to the smallest source size parameter, $R=1\ \mathrm{fm}$. In the case of model A, at threshold, we have $C_{D^{*}\pi}^{(1/2)}(k=0)>1$, because of the attractive character of this channel and the negative scattering length. In the sequence, as $k$ increases a moderate minimum and a bump are found in the region $220\lesssim k\lesssim 360\ \mathrm{MeV}$. Interestingly, these effects reflect essentially the behavior of the $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$ amplitude, since the other contributions $T_{D^{*}\pi,D\rho}^{(1/2)}$ and $T_{D^{*}\pi,\bar{K}D_{s}^{*}}$ are negligible, as shown in Fig. 9  [^9] This finding is in agreement with the result on the $D^{*}\pi$ amplitude discussed in section II.3.1 and II.3.2. An equivalent effect is also found in the femtoscopic analysis of the coupled-channel $N\Xi$ and $\Lambda\Lambda$ interactions Kamiya et al. (2022).. In this sense, the minimum (bump) at $k\gtrsim 250\ \mathrm{MeV}$ ($k\simeq 340\ \mathrm{MeV}$) is associated to the broad peak (dip) in $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$ at $\sqrt{s}\sim 2304\ \mathrm{MeV}$ ($\sqrt{s}\sim 2405\ \mathrm{MeV}$). Thus, according to our model, the CF encodes the manifestation of the interference between the poles present in $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$ (see discussions in section II.3.1). However, these effects are no longer prominent for larger values of the source size parameter. Also, a cusp at $k\simeq 518\ \mathrm{MeV}$ is seen and comes from the effect of the $\bar{K}D_{s}^{*}$ threshold. Notably, when compared to the single-channel LL results, these CFs have a similar qualitative behavior only near the threshold, having values larger than one. The $C_{LL}$ does not acquire a minimum from the peak in $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$, possibly because of its large width (as argued in Sec. III.2).
@@
Figure 9: CFs with $I=1/2$ for the $D^{*}\pi$ (left panel) and $D\rho$ (right panel) as functions of their relative momentum $k$ considering the single contribution of the elastic channel (i.e. only $i=j$ of the sum in Eq. (23)), and the total CF.
@@ III.3.2 CFs in physical basis
We remark that the CFs presented so far, for the relevant channels $D^{*}\pi$ and $D\rho$, are in the isospin basis. Therefore, in order to provide measurable CFs, we need to express them on the particle basis. In the case of $D^{*}\pi$ (the case of $D\rho$ is completely analogous), we consider the isospin-doublet of the vector charmed meson and isospin-triplet of the pion as $D\equiv(|D^{*+}\rangle,-|D^{*0}\rangle)$ and $\pi\equiv(-|\pi^{+}\rangle,|\pi^{0}\rangle,|\pi^{-}\rangle)$, respectively. Then, for $D^{*}\pi$-states with $I_{3}=+1/2$, the particle basis is given by ${|D^{*0}\pi^{+}\rangle,|D^{*+}\pi^{0}\rangle}$, which is related to the isospin basis through (denoting states as $|D\pi,I\rangle$)
@@
We show the CFs for the $D^{*}\pi$ and $D\rho$ states in the particle basis defined in Eq. (34), in Fig. 10, for both models A and B. The CFs $C_{D^{*0}\pi^{-}}(k)$ and $C_{D^{0}\rho^{-}}(k)$ are also plotted; since these states have $I_{3}=-3/2$, their corresponding CFs are naturally equal to $C_{i}^{(3/2)}(k)$. It can be seen that for model A the features of the $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$ amplitude are more notable in the channel $D^{*0}\pi^{+}$, because of the bigger weight of the $I=1/2$ channel in its wave function. In contrast, for model B there is no sizable difference among the channels $D^{*0}\pi^{+},D^{*+}\pi^{0}$ and $D^{*0}\pi^{-}$, due to the similarity among $C_{D^{*}\pi}^{(1/2)}(k)$ and $C_{D^{*}\pi}^{(3/2)}(k)$. Going to the scenario of $D\rho$, the difference coming from the isospin weights produces $C_{D^{+}\rho^{0}}(k)$ closer to one at threshold than $C_{D^{0}\rho^{+}}(k)$. Hence, one can conclude that the $D^{*0}\pi^{+}$ and $D^{0}\rho^{+}$ channels are more appropriate to test both models.
In the end, we stress the main conclusion of this study, namely: our findings suggest that $C_{D^{*}\pi}(k)$ and $C_{D\rho}(k)$ might encode sufficiently identifiable signatures of the $D_{1}(2430)^{0}$ and $D_{1}(2420)$ states when smaller sources are considered. It should be emphasized that if the femtoscopic analysis for the mentioned channels is done and the measured genuine CFs have similar behavior to those obtained here, then it is possible to say that this work provides a framework compatible with the existence of both broad and narrow states. In this sense, it would be interesting to confront these CFs with data collected in future high precision experiments.
@@ IV Conclusions
The main conclusions of the discussions presented in this work can be summarized as follows.
  • The information on the properties of the two lightest $D_{1}$ states comes mostly from fits made to the experimental data on the $D^{*}\pi$ invariant mass, which gets contribution from several charm states. Such a procedure attributes similar masses but very different widths to the two of them, which indicates that the two states must have something different in their nature.
  • There exists information on the scattering length, determined in the lattice QCD calculations, for the $D^{*}\pi$ and $D\pi$ channels. Both values are very similar. Several model calculations constrain their parameters using the former information and determine the $D\pi$ scattering length in the infinite volume by using physical masses. It can be argued that the scattering lengths for $D\pi$ and $D^{*}\pi$ can be similar, which provides another valuable information on the study of the $D^{*}\pi$ system.
  • A value for the $D\pi$ scattering length is also available from heavy ion collisions but it does not agree with those mentioned in the previous point.
  • With the purpose of understanding the properties of the two enigmatic $D_{1}$ states, we consider a model where different meson-meson interactions and a bare quark model pole constitute the lowest order amplitudes. Such amplitudes are used as kernels to solve the Bethe-Salpeter equation in a coupled channel approach. Consequently, a narrow pole is found to get generated by the hadron dynamics and is related to $D_{1}(2420)$. A broader pole is also found, whose properties match those of $D_{1}(2430)$ when both hadron dynamics and a bare quark-model pole are considered.
@@
This work is partly supported by the Brazilian agencies CNPq (L.M.A.: Grant Numbers 309950/2020-1, 400215/2022-5, 200567/2022-5), FAPESP (K.P.K.: Grant Number 2022/08347-9; A. M. T.: Grant number 2023/01182-7); FAPESB (L.M.A.: Grant Number INT0007/2016); and CNPq/FAPERJ under the Project INCT-Física Nuclear e Aplicações (Contract No. 464898/2014-5).

### v2

@@ III.2 Lednicky-Lyuboshits approximation
To shed some light on the interpretation of the CFs, it can be instructive to review the Lednicky-Lyuboshits (LL) model, which is based on replacing the full wave function for a single channel by its nonrelativistic, asymptotic ($r\to\infty$) form, corresponding to the superposition of plane and converging spherical waves Lednicky and Lyuboshits (1981). In particular, we benefit from the discussion presented in the Appendix of Ref. Albaladejo et al. (2023) and Secs. V.B and V.C of Ref. Kamiya et al. (2022), which have some of their fundamental aspects reproduced here.
Proceeding ahead, the consideration of the LL approximation, together with a Gaussian source, and using the relationship between the standard quantum mechanics amplitude $f(k)$ and the scattering matrix $T$, i.e., $f(k)=-T/(8\pi\sqrt{s})$, allow us to write the single-channel CF as Fabbietti et al. (2021); Lednicky and Lyuboshits (1981); Kamiya et al. (2022); Albaladejo et al. (2023)
@@
Thus, considering an attractive interaction generated by the strong force, the CF given in Eqs. (31) and (32) behaves as follows. Near the threshold, for negative values of the scattering length (which means an unbound scenario for the system) the CF acquires (i) a strong enhancement when $|a|\gg R$ (i.e., smaller source), (ii) a moderate enhancement when $|a|\sim R$, and (iii) a value $C_{LL}(0)\gtrsim 1$ when $|a|\ll R$ (larger source). On the other hand, in the situation of $a>0$, corresponding to an attractive interaction which could generate a bound or quasibound state, in the low-momentum limit, the CF achieves (i) a strong enhancement for smaller sources $(a\gg R)$, (ii) a value very close to one when $a/R\sim 4/\sqrt{\pi}\sim 2.3$, (iii) its minimum value ($\simeq 0.4$) when $a/R\sim 2/\sqrt{\pi}\simeq 1.1$, (iv) a moderate dip when $a<R$; and (v) a value $C_{LL}(0)\lesssim 1$ for larger sources ($a\ll R$). This behavior is summarized in Fig. 7 (left panel), where one can see that the enhancement of CF at a given value of $R$ is not conclusive concerning the formation of a bound or quasibound state. Notwithstanding this, as a consequence of the dependence of the CF on $R$ and $a$, one can infer the existence of a bound or quasibound state when, near the threshold, the CF moves from an enhancement to a dip at $a\simeq R$ as $R$ increases. In this sense, experimental analyses of the CF in systems with different sizes, for instance, $pp,pA$, and $AA$ collisions, deserve special attention.
@@
Then, considering that a resonance present at $k_{R}$ generates $\delta(k_{R})=\pi/2$, which when used in Eq. (33) yields $C_{LL}(x_{R})=1+\frac{e^{-4x_{R}^{2}}}{2x_{R}^{2}};\ x_{R}=k_{R}R$. Accordingly, the CF ends up having the following properties (i) $C_{LL}^{\prime}(x_{R})<0$, (ii) $C_{LL}<1$ for $x\gtrsim x_{R}$, and (iii) $C_{LL}\simeq 1$ for large $x_{R}$. To be more didactic, from the analytical expression for the Breit-Wigner-like phase shift, $\delta(E)=\Gamma(E)/(2E_{R}-2E)$ ($\Gamma$ being the width and $E_{R}$ the energy at $k_{R}$), one gets $\cot{\delta(k)}\sim-(x^{2}-x_{R}^{2})/x_{\Gamma}^{2}$, with $x_{\Gamma}^{2}=\mu\Gamma R^{2}$ ($\mu$ being the reduced mass of the two particles in the channel). Thus, the use of this last expression of $\cot{\delta(k)}$ in Eq. (33) can engender in $C_{LL}$ (i) a maximum at $k\lesssim k_{R}$ and a pronounced minimum at $k\gtrsim k_{R}$ for small $x_{\Gamma}$, (ii) a weakened minimum at $k\gtrsim k_{R}$ for intermediate values of $x_{\Gamma}$, and (iii) an almost plateaulike appearance at $k\lesssim k_{R}$ for large $x_{\Gamma}$. This behavior is summarized in Fig. 7 (right panel) in accordance with that in Ref. Albaladejo et al. (2023).
@@ III.3.1 CFs in isospin basis
Figure 8: CFs for the $D^{*}\pi$ (top panels) and $D\rho$ (bottom panels) channels as functions of their relative momentum $k$, taking different values of the source size. The left, center, and right panels show the results obtained, respectively, with $I=1/2$ considering the model A, with $I=1/2$ considering model B, and with $I=3/2$. Recall that the interaction in the isospin 3/2 configuration is repulsive and weaker as compared to the case of isospin 1/2. Thus, no states are formed in the isospin 3/2 case and the amplitudes are obtained from the same model [using Eq. (2)].
@@
In Fig. 8 we plot the results of the correlation functions for the most relevant channels $D^{*}\pi$ and $D\rho$ as functions of their c.m. momentum $k$, for different values of the range parameter $R$ of the source, considering the following scenarios: $I=1/2$ with model A, $I=1/2$ with model B and $I=3/2$. For the sake of comparison and to reach a more profound comprehension concerning our findings, the results with the single-channel LL approximation are also included.
@@
First, one can notice the distinct behavior of the $C_{D^{*}\pi}^{(1/2)}(k)$ when models A or B are employed ( see left and center panels in the upper row of Fig. 8). To start with the discussions, let us focus on the curves related to the smallest source size parameter, $R=1\ \mathrm{fm}$. In the case of model A, at threshold, we have $C_{D^{*}\pi}^{(1/2)}(k=0)>1$, because of the attractive character of this channel and the negative scattering length. In the sequence, as $k$ increases a moderate minimum and a bump are found in the region $220\lesssim k\lesssim 360\ \mathrm{MeV}$. Interestingly, these effects reflect essentially the behavior of the $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$ amplitude, since the other contributions $T_{D^{*}\pi,D\rho}^{(1/2)}$ and $T_{D^{*}\pi,\bar{K}D_{s}^{*}}$ are negligible, as shown in Fig. 9  [^10] This finding is in agreement with the result on the $D^{*}\pi$ amplitude discussed in section II.3.1 and II.3.2. An equivalent effect is also found in the femtoscopic analysis of the coupled-channel $N\Xi$ and $\Lambda\Lambda$ interactions Kamiya et al. (2022).. In this sense, the minimum (bump) at $k\gtrsim 250\ \mathrm{MeV}$ ($k\simeq 340\ \mathrm{MeV}$) is associated to the broad peak (dip) in $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$ at $\sqrt{s}\sim 2304\ \mathrm{MeV}$ ($\sqrt{s}\sim 2405\ \mathrm{MeV}$). Thus, according to our model, the CF encodes the manifestation of the interference between the poles present in $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$ (see discussions in Sec. II.3.1). However, these effects are no longer prominent for larger values of the source size parameter. Also, a cusp at $k\simeq 518\ \mathrm{MeV}$ is seen and comes from the effect of the $\bar{K}D_{s}^{*}$ threshold. Notably, when compared to the single-channel LL results, these CFs have a similar qualitative behavior only near the threshold, having values larger than one. The $C_{LL}$ does not acquire a minimum from the peak in $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$, possibly because of its large width (as argued in Sec. III.2).
@@
Figure 9: CFs with $I=1/2$ for the $D^{*}\pi$ (left panel) and $D\rho$ (right panel) as functions of their relative momentum $k$ considering the single contribution of the elastic channel [i.e., only $i=j$ of the sum in Eq. (25)], and the total CF.
@@ III.3.2 CFs in physical basis
We remark that the CFs presented so far, for the relevant channels $D^{*}\pi$ and $D\rho$, are in the isospin basis. Therefore, in order to provide measurable CFs, we need to express them on the particle basis. In the case of $D^{*}\pi$ (the case of $D\rho$ is completely analogous), we consider the isospin doublet of the vector charmed meson and isospin triplet of the pion as $D\equiv(|D^{*+}\rangle,-|D^{*0}\rangle)$ and $\pi\equiv(-|\pi^{+}\rangle,|\pi^{0}\rangle,|\pi^{-}\rangle)$, respectively. Then, for $D^{*}\pi$ states with $I_{3}=+1/2$, the particle basis is given by ${|D^{*0}\pi^{+}\rangle,|D^{*+}\pi^{0}\rangle}$, which is related to the isospin basis through (denoting states as $|D\pi,I\rangle$)
@@
We show the CFs for the $D^{*}\pi$ and $D\rho$ states in the particle basis defined in Eq. (36), in Fig. 10, for both models A and B. The CFs $C_{D^{*0}\pi^{-}}(k)$ and $C_{D^{0}\rho^{-}}(k)$ are also plotted; since these states have $I_{3}=-3/2$, their corresponding CFs are naturally equal to $C_{i}^{(3/2)}(k)$. It can be seen that for model A the features of the $T_{D^{*}\pi,D^{*}\pi}^{(1/2)}$ amplitude are more notable in the channel $D^{*0}\pi^{+}$, because of the bigger weight of the $I=1/2$ channel in its wave function. In contrast, for model B there is no sizable difference among the channels $D^{*0}\pi^{+},D^{*+}\pi^{0}$, and $D^{*0}\pi^{-}$, due to the similarity among $C_{D^{*}\pi}^{(1/2)}(k)$ and $C_{D^{*}\pi}^{(3/2)}(k)$. Going to the scenario of $D\rho$, the difference coming from the isospin weights produces $C_{D^{+}\rho^{0}}(k)$ closer to one at threshold than $C_{D^{0}\rho^{+}}(k)$. Hence, one can conclude that the $D^{*0}\pi^{+}$ and $D^{0}\rho^{+}$ channels are more appropriate to test both models. It is worth mentioning that correlation functions for $D^{*+}\pi^{+}$ and $D^{*+}\pi^{-}$ have been calculated in Ref. Torres-Rincon et al. (2023), although with very different poles present in the amplitudes, those in line with Refs. Abreu et al. (2019); Du et al. (2018).
In the end, we stress the main conclusion of this study, namely: our findings suggest that $C_{D^{*}\pi}(k)$ and $C_{D\rho}(k)$ might encode sufficiently identifiable signatures of the $D_{1}(2430)^{0}$ and $D_{1}(2420)$ states when smaller sources are considered. It should be emphasized that if the femtoscopic analysis for the mentioned channels is done and the measured genuine CFs have similar behavior to those obtained here, then it is possible to say that this work provides a framework compatible with the existence of both broad and narrow states. Determining information on the $D^{*}\pi$ channel, where both mesons are not electrically charged, from smaller sources, like proton-proton collisions, can be useful in settling the value of the $D^{*}\pi$ scattering length. It should be possible to determine data on $D^{*0}\pi^{+}$ where $D^{*0}$ is reconstructed from $D^{+}\pi^{-}$. In this sense, it would be interesting to confront these CFs with data collected in future high-precision experiments.
@@ IV Conclusions
The main conclusions of the discussions presented in this work can be summarized as follows:
  • The information on the properties of the two lightest $D_{1}$ states comes mostly from fits made to the experimental data on the $D^{*}\pi$ invariant mass, which gets contribution from several charm states. Such a procedure attributes similar masses but very different widths to the two of them. Although two states with similar masses are expected from the quark model, and once the masses are assumed one could explain their different widths on the basis of heavy quark symmetry, it seems not trivial to simultaneously describe both (masses and widths) from the same dynamics. Different models lead to different results and considerations of hadronic loops or similar mechanisms of mixing between $11^{1}P$ 1 and $11^{3}P$ 1 quark model seems to work better.
  • There exists information on the scattering length, determined in the lattice QCD calculations, for the $D^{*}\pi$ and $D\pi$ channels. Both values are very similar. Several model calculations constrain their parameters using the former information and determine the $D\pi$ scattering length in the infinite volume by using physical masses. It can be argued that the scattering lengths for $D\pi$ and $D^{*}\pi$ can be similar, which provides valuable information on the study of the $D^{*}\pi$ system.
  • A value for the $D^{*}\pi$ scattering length is also available from heavy ion collisions but it does not agree with those mentioned in the previous point.
  • With the purpose of understanding the properties of the two $D_{1}$ states and finding alternative ways to extract information on the $D^{*}\pi$ scattering length, we consider a model where different meson-meson interactions and a bare quark model pole constitute the lowest order amplitudes. Such amplitudes are used as kernels to solve the Bethe-Salpeter equation in a coupled channel approach. Consequently, a narrow pole is found to get generated by the hadron dynamics and is related to $D_{1}(2420)$. A broader pole is also found, whose properties match those of $D_{1}(2430)$ when both hadron dynamics and a bare quark-model pole are considered.
@@ V Acknowledgements
This work is partly supported by the Brazilian agencies CNPq (L.M.A.: Grants No. 309950/2020-1, No. 400215/2022-5, No. 200567/2022-5, and No. 308299/2023-0, K.P.K.: Grants No. 407437/2023-1 and No. 306461/2023-4, A.M.T: Grant No. 304510/2023-8), FAPESP (K.P.K.: Grant No. 2022/08347-9; A. M. T.: Grant No. 2023/01182-7); and CNPq/FAPERJ under the Project INCT-Física Nuclear e Aplicações (Contract No. 464898/2014-5).